flux, what is the basketball flux through the hoop in meters squared per second? Assume the rim is a circle with a radius of 0.23 meters. Hint: recall the equation for the area of a circle a. 5.2 x 105 basketballs m2 s1 b. 5.2 x 105 basketballs m-2 s-1 c. 6.3 x 102 basketballs m2 s1 6.3x 102 basketballs m2 s+
Answer #1

C) 6.3 * 10^-2 is the ans.

30 shots in 48mins =2880 seconds

1 shot in 2880/30= 96 seconds

area of circle = 3.14*0.23*0.23=0.166 m2

flux=(1/0.166*96)=6.3 * 10^-2 m^-2s^-1